OpenStudy (anonymous):

Find the distance between (1, -2) and (-4, 3).

pooja195 (pooja195):

\[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] Use the distance formula

OpenStudy (anonymous):

Ok what next

pooja195 (pooja195):

\[\huge d=\sqrt{(-4-1)^2+(3-(-2)^2}\] \[\huge d=\sqrt{(-4-1)^2+(3+2)^2}\]

pooja195 (pooja195):

Lets start with the first one -4-1=?

pooja195 (pooja195):

If you need help let me know :)